OpenStudy (anonymous):

How to find the horizontal asymptote for Cos(x)+1 or any other simple trigonometric function ?

OpenStudy (dumbcow):

i assume you are referring to the max/min of sin and cosine functions A = amplitude k = vertical shift max = A +k min = -A +k Ex: cos x + 1 A = 1 k = 1 max = 2, min = 0

OpenStudy (anonymous):

Thanks alot